Skip to main contentSkip to solution

A farmer has a tubewell which works by connecting two switches S1S_1 and S2S_2. The probability that S1S_1 works is 9/10 and the probability that S2S_2 works is 8/10. The farmer connects them in series in 30% cases and in parallel in 70% cases. If it is known that the tubewell was working, what is the probability that it was a series connection.

Solution

✅ Correct Option: 2

In series the tubewell works with probability 0.9×0.8=0.720.9\times 0.8 = 0.72; in parallel with probability 1−0.1×0.2=0.981 - 0.1\times 0.2 = 0.98.

Total probability of working =0.3(0.72)+0.7(0.98)=0.216+0.686=0.902= 0.3(0.72) + 0.7(0.98) = 0.216 + 0.686 = 0.902.

By Bayes' theorem, the required probability is 0.2160.902=108451\dfrac{0.216}{0.902} = \dfrac{108}{451}.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question